# Uncertainty Quantification on Graph Learning: A Survey

### v2

@@ IV-C1 Gaussian Process Overview
Gaussian Processes (GPs) are a widely-used class of non-parametric Bayesian models that have been widely used for uncertainty quantification in machine learning [71]. A GP is defined as a collection of random variables, any finite number of which have a joint Gaussian distribution. In the context of function approximation, a GP can be used as a prior distribution over the space of functions, denoted as:
@@
where $m(\mathbf{x})$ is the mean function and $k(\mathbf{x},\mathbf{x^{\prime}})$ is the covariance function or kernel. The mean function represents the expected value of the function at any input point, while the covariance function captures the similarity between pairs of input points and determines the smoothness of the functions sampled from the GP.
Given a set of observed data $\mathcal{D}={(\mathbf{x}_{i},y_{i})}_{i=1}^{n}$, where $\mathbf{x}_{i}$ are the input points and $y_{i}$ are the corresponding target values, the posterior distribution is a GP with updated mean and covariance functions:
@@
where $\mathbf{k}(\mathbf{x})=[k(\mathbf{x},\mathbf{x}_{1}),\dots,k(\mathbf{x},\mathbf{x}_{n})]^{T}$, $\mathbf{K}$ is the covariance matrix (or kernel matrix) with entries $\mathbf{K}_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$, $\mathbf{y}=[y_{1},\dots,y_{n}]^{T}$, $\mathbf{m}=[m(\mathbf{x}_{1}),\dots,m(\mathbf{x}_{n})]^{T}$, and $\sigma^{2}_{n}$ is the noise variance.
The posterior distribution allows for uncertainty quantification in predictions. For a new input point $\mathbf{x}_{*}$, the predictive distribution is given by:
$p(f_{*}|\mathbf{x}_{*},\mathcal{D})=\mathcal{N}(f_{*}|\mu(\mathbf{x}_{*}),\Sigma(\mathbf{x}_{*},\mathbf{x}_{*})).$ (25)
@@
The predictive mean $\mu(\mathbf{x}_{*})$ provides the expected value of the function at $\mathbf{x}_{*}$, while the predictive variance $\Sigma(\mathbf{x}_{*},\mathbf{x}_{*})$ quantifies the uncertainty associated with the prediction.
@@
One limitation of standard GPs is their computational complexity, which scales cubically with the number of training points due to the inversion of the covariance matrix. Various approximation techniques have been proposed to overcome this limitation [72, 73, 74], with the added benefit of enabling efficient UQ in GPs. Recent advancements have also explored the integration of GPs with deep learning architectures to leverage the strengths of both approaches [75, 76]. This integration enhances the ability to quantify uncertainty in complex and hierarchical data representations.
IV-C2 Gaussian Process for Graphs
@@
Applying GPs to graph-structured data has emerged as an important research direction in recent years [77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. This approach combines the powerful UQ capabilities of GPs with the complex structures of graph data, offering new possibilities for uncertainty estimation in graph learning tasks. This section focuses on works that are closely related to UQ in graph learning, while also introducing other relevant GP methods.
@@
Closely Related to UQ. [77] were among the first to propose a Graph Gaussian Process (GGP) based semi-supervised learning method. This method can achieve good generalization performance even with limited labeled data, bridging the gap between simple models and data-intensive graph neural networks. GGP defines a GP prior over node embeddings, using a kernel function that can be selected from existing well-studied options. The model combines the GP prior with a robust-max likelihood [87] to handle multi-class problems. Since the posterior distribution is not analytically tractable, a variational inference-based approximation method with inducing points is used, similar to the inter-domain inference algorithm [88].
[78] proposed Uncertainty aware Graph Gaussian Process (UaGGP), an extension of the GGP model. UaGGP addresses the uncertainty in graph structures that may arise in semi-supervised learning problems on graphs, where edges can connect data points with different labels, leading to uncertainty in predictions. UaGGP uses GP to guide the model learning by incorporating prediction uncertainty and label smoothness regularization. For any two nodes $v_{i}$ and $v_{j}$ in the graph, UaGGP defines an uncertainty-aware distance measure $d(v_{i},v_{j})$ to constrain the proximity of their prediction results. When the model is not confident in the inference of $v_{i}$ or $v_{j}$ or both of them, the distance $d(v_{i},v_{j})$ should not be small.
To detect potential failures of GNNs on OOD samples in the task of energy and force predictions on molecules, [84] proposed a method called Localized Neural Kernel (LNK) to address the uncertainty estimation problem in molecular force fields. LNK leverages GNNs as encoders within a Deep Kernel Learning framework and uses Sparse Variational Gaussian Processes (SVGP) to mitigate the computational burden associated with GPs. By directly fitting the GP on atomic embeddings, this method ensures the locality of uncertainty in predictions, particularly effective in modeling molecular interactions.
Other GP Methods for Graph Learning. [79] introduced GPGC, a GP regression model using GCNs that enhances expressive power by focusing on 1-hop neighborhood nodes with equal weights. [80] incorporated graph structures into the GP kernel using spectral graph wavelets, adopting a multi-scale approach to capture various levels of smoothness and adapt to different frequencies. Similarly, [81] proposed the GPG model, integrating the graph Laplacian into the GP covariance matrix to enhance spectral regularization of target signals in node regression tasks.
To address scalability in large-scale graph data, [82] developed Stochastic Deep Gaussian Processes over Graphs (DGPG), which leverages deep GPs to model complex mappings efficiently. In the context of graph classification, [83] introduced two models, FT-GP and WT-GP, utilizing graph Fourier and wavelet transforms to extract spectral features suitable for the GP framework, capturing multi-scale and localized patterns in graphs.
For link prediction, [85] proposed a graph convolutional variational GP model that combines graph convolution operations with GP to effectively utilize graph structures and node features. Lastly, [86] introduced the Multi-Relational GGPN, which employs GPs to model distributed embedding functions, allowing for the learning of a family of embedding functions and enhancing the model’s flexibility.
@@ V Methods for Uncertainty Handling
V-A Out of Distribution
Out-of-distribution (OOD) refers to instances that significantly differ from a model’s training data. It is especially crucial for models in real-world applications that face unpredictable and uncertain inputs, in stark contrast to the controlled and uniform datasets used for training. When a model encounters OOD data, its epistemic uncertainty typically increases. OOD data typically lead to the model’s underperformance or unpredictable behavior because of its great difference from the training data, resulting in high prediction uncertainties. To address this problem, a surge of research has focused on out-of-distribution (OOD) generalization. Strategies to mitigate uncertainty issues posed by OOD generalization can be broadly classified into several categories [89], including distributionally robust optimization (DRO) [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103], adversarial training (AT) [104, 105, 106], and self-supervised learning (SSL) [107, 108, 109, 110]. Most of them propose novel frameworks, network architectures, and objective functions to improve the model performance of OOD data during testing. However, this survey focuses on the measurement, calibration, and conformity of predictive uncertainty; thus, we will emphasize the integration of OOD and uncertainty studies, especially how to measure and estimate the disagreement between training and test data.
V-A1 Distributionally Robust Optimization

### v3

@@ IV-B1 Gaussian Process Overview
Gaussian Processes (GPs) are a popular class of non-parametric Bayesian models that have been widely used for uncertainty quantification in machine learning [139]. A GP is defined as a collection of random variables, any finite number of which have a joint Gaussian distribution. In the context of function approximation, a GP can be used as a prior distribution over the space of functions, denoted as:
@@
where the mean function $m(\mathbf{x})$ calculates the expectation of the function at input $\mathbf{x}$. The covariance function $k(\mathbf{x},\mathbf{x^{\prime}})$ captures the similarity between pairs of inputs and determines the smoothness of the functions sampled from the GP.
Given a set of observed data $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, where $\mathbf{x}_{i}$ are the input points and $y_{i}$ are the corresponding target values, the posterior distribution is a GP with updated mean and covariance functions:
@@
where $\mathbf{k}(\mathbf{x})=[k(\mathbf{x},\mathbf{x}_{1}),\dots,k(\mathbf{x},\mathbf{x}_{n})]^{T}$, $\mathbf{K}$ is the covariance matrix (or kernel matrix) with entries $\mathbf{K}_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$, $\mathbf{y}=[y_{1},\dots,y_{n}]^{T}$, $\mathbf{m}=[m(\mathbf{x}_{1}),\dots,m(\mathbf{x}_{n})]^{T}$, and $\sigma^{2}_{n}$ is the noise variance. The posterior distribution allows for uncertainty quantification in predictions. For a new input point $\mathbf{x}^{*}$, the predictive distribution is given by:
$p(f^{*}|\mathbf{x}^{*},\mathcal{D})=\mathcal{N}(f^{*}|\mu(\mathbf{x}^{*}),\Sigma(\mathbf{x}^{*},\mathbf{x}^{*})).$ (23)
@@
The predictive mean $\mu(\mathbf{x}^{*})$ provides the expected value of the function at $\mathbf{x}^{*}$, while the predictive variance $\Sigma(\mathbf{x}^{*},\mathbf{x}^{*})$ quantifies the uncertainty associated with the prediction.
@@
To address the computational cost of standard GPs, efficient approximations have been developed to preserve Bayesian uncertainty quantification at reduced cost [140, 107, 141]. Recent advancements have also explored the integration of GPs with deep learning architectures to leverage the strengths of both approaches [142, 143]. This integration enhances the ability to quantify uncertainty in complex data representations.
IV-B2 Gaussian Process for Graphs
@@
Applying GPs to graph-structured data has emerged as an important research direction in recent years [64, 65, 122, 123, 124, 125, 126, 66, 127, 128]. This approach combines the powerful UQ capabilities of GPs with the complex structures of graph data, offering new possibilities for uncertainty estimation in graph learning tasks. The following discusses works closely related to UQ in graph learning, along with other relevant GP methods.
@@
[64] were among the first to propose a Graph Gaussian Process (GGP) based semi-supervised learning method. This method can achieve good generalization performance even with limited labeled data, bridging the gap between simple models and data-intensive graph neural networks. GGP defines a GP prior over node embeddings, using a kernel function that can be selected from existing well-studied options. The model combines the GP prior with a robust-max likelihood [144] to handle multi-class problems. As an extension of the GGP model, UaGGP [65] addresses the uncertainty in graph structures that may arise in semi-supervised learning problems on graphs, where edges can connect data points with different labels, leading to uncertainty in predictions. UaGGP uses GP to guide the model learning by incorporating prediction uncertainty and label smoothness regularization. For any two nodes $v_{i}$ and $v_{j}$, UaGGP defines an uncertainty-aware distance measure $d(v_{i},v_{j})$ to constrain the proximity of the predictions at those nodes. This mechanism manifests uncertainty quantification by ensuring that $d(v_{i},v_{j})$ increases whenever the model’s confidence in one or both node predictions is low.
Subsequent extensions address scalability and specialized tasks: DGPG [125] proposes stochastic deep GPs for scalable node regression, while LNK [66] combines GNN-based encoders with sparse variational GPs for uncertainty estimation in molecular energy and force predictions. FT-GP and WT-GP [127] employ Fourier and wavelet transforms to capture multiscale and localized spectral information for graph classification. Similarly, GPG [124] integrates the graph Laplacian into the GP covariance matrix to enhance spectral regularization of target signals. GGPN [128] uses GPs to learn distributed embedding functions, enhancing flexibility in representing diverse graph structures.
IV-C OOD Epistemic Uncertainty Representation
Enhancing the generalization is an effective strategy for improving models’ performance under uncertainty. It enables models to make accurate predictions for “unexpected” and previously unseen samples. These outlier samples, which significantly differ from the training (in-distribution, ID) data, are commonly referred to as out-of-distribution (OOD) samples. In this subsection, we survey various approaches to model perturbations and represent outliers.
IV-C1 Dataset-level uncertainty estimation
One prominent approach to uncertainty representation is through dataset-level uncertainty estimation [145], where the model accounts for distributional shifts by defining an uncertainty set over data distributions. In this way, OOD or unseen data are modeled as elements outside a specified neighborhood of the empirical training distribution.
Formally, the uncertainty set is defined as $\mathcal{P}:=\{P|D(P,P_{0})\leq\rho\}$, where $P_{0}$ is the observed data distribution, and $\rho$ denotes a radius distance. The distance between two distributions, $P$ and $P_{0}$, is measured by a divergence metric $D(\cdot,\cdot)$. We list the most widely used choices of $D$.
[…]
  • KL-divergence, one special case of $f$-divergence, is used as a divergence measure [150, 67, 151, 152]. However, the possible limitations of metrics based on $f$-divergence, including KL-divergence, are that they may not be comprehensive enough to include some relevant distributions, or they may fail to precisely measure the divergence for extreme distributions [153].
[…]
  • Besides that, authors in [160] study uncertainty sets defined by maximum mean discrepancy, and Erdoğan and Garud [161] define the uncertainty sets using the Prohorov metric. Formally, $D_{Pro}(P,P_{0}):=\inf{\{\epsilon>0|P(X)\leq P_{0}(X^{\epsilon})+\epsilon\text{, }P_{0}(X)\leq P(X^{\epsilon})+\epsilon\}}$, where $X^{\epsilon}$ is the $\epsilon$-neighborhood of $X$.
  • Moreover, Wu et al. [162] define the graph-level distance by a novel graph subtree discrepancy based on Weisfeiler-Lehman Subtree Kernel. Fang et al. [152] define a graph-level heterophily distribution.
These dataset-level uncertainty estimation provides a principle way to incorporate distributional robustness into learning objectives, enabling models to represent and generalize better under real-world data variability and shift.
IV-C2 Sample-level uncertainty representation
Sample-level uncertainty refers to the uncertainty in the model’s prediction for one sample. For example, Qian et al. [145] measure it by perturbing an input sample and observing the variance in its predicted probabilities. This type of uncertainty quantifies how much the model’s prediction can change when the input is slightly modified. Adversarial samples are a prime illustration of such uncertainty.
By adopting adversarial attack methods such as FGSM [163], PGD [164], or CW attacks [165], one can generate adversarial samples that manipulate the model’s predictions by very limited perturbations. These adversarial samples essentially highlight regions where the model is uncertain and highly sensitive to small changes, and thus serve as a practical proxy to quantify sample-level uncertainty, with regions of the input space that produce significant shifts in model predictions being areas of high uncertainty. Formally, adversarial attacks can be defined in two alternative ways:
$\epsilon^{\ast}=\arg\max_{\|\epsilon\|\leq\rho}\mathcal{L}_{0}(\theta,x+\epsilon,y),$ (24)
or
$\epsilon^{\ast}=\arg\min_{\epsilon}\|\epsilon\|\;\;\textnormal{s.t. }f(x+\epsilon)\neq f(x),$ (25)
where $(x,y)$ denotes a sample with its corresponding label, and the sample $x$ is manipulated by an adversarial perturbation $\epsilon$, constrained by a predefined small budget $\rho$. $x+\epsilon^{\ast}$ is the adversarial counterpart of original input $x$. The general loss function $\mathcal{L}_{0}$, such as cross-entropy loss, quantifies the model’s performance on the specific task.
IV-C3 Manually designed uncertainty sets
OOD epistemic uncertainty arises from the lack of knowledge about certain inputs. When imposing a formal probability distribution on inputs or model parameters is hard, we can approximate this uncertainty by generating controlled perturbations of each input. These perturbations are carefully designed to preserve the core semantic or structural content while introducing realistic variations. By observing how the model’s predictions vary across these perturbations, we obtain a proxy measure of epistemic uncertainty.
Specifically, uncertainty sets are constructed by applying controlled perturbations while retain key structural or attributional information. By adopting controlled perturbations, input graphs are augmented in a manually controllable way. Structural augmentations include node/edge deletion or addition [91], and subgraph sampling via random walks to preserve local topology [74]. At the attribute level, node and edge features can be masked or corrupted to mimic missing or noisy information. Some works combine both structural and attribute perturbations to generate diverse graph augmentations [93].
These manually designed perturbations offer interpretable and task-relevant ways to simulate uncertainty, particularly when distributional assumptions are hard to formalize.
IV-D Stochastic Differential Equations
Stochastic Differential Equations (SDEs) offer a dynamic systems perspective for modeling uncertainty in neural networks. Instead of viewing network transformations as deterministic functions, SDE-based approaches model the evolution of hidden states or network parameters as stochastic processes. This continuous-time formulation provides a natural way to incorporate and quantify uncertainty throughout the network.
IV-D1 SDEs Overview
The core idea behind using SDEs for UQ in neural networks is to represent the mapping from input to output, or the learning dynamics of parameters, as the solution to an SDE. A general form of an SDE for a state process $\mathbf{x}_{t}$ is:
$d\mathbf{x}_{t}=f(\mathbf{x}_{t},t)dt+g(\mathbf{x}_{t},t)dW_{t}$ (26)
Here, $f(\mathbf{x}_{t},t)$ is the drift term, dictating the deterministic part of the evolution, and $g(\mathbf{x}_{t},t)$ is the diffusion term, scaling the stochastic part driven by a Wiener process (Brownian motion) $dW_{t}$. Both drift and diffusion functions can be parameterized by neural networks.
SDE-Net [166] instantiates the general SDE in Eq. (26) by using two neural networks. The first network (“drift net”) models the deterministic term. The second network (“diffusion net”) governs the stochastic term. The diffusion net induces a distribution over the final hidden state $\mathbf{x}_{T}$. Its variance, $\mathrm{Var}(\mathbf{x}_{T})$, quantifies epistemic uncertainty. Aleatoric uncertainty is then obtained from the predictive distribution $p(y\mid\mathbf{x}_{T})$, measured via expected predictive entropy for classification and predictive variance for regression.
Building on similar principles, other works have explored SDEs for UQ in various contexts. For instance, SDEs have been employed to conceptualize infinitely deep Bayesian neural networks, allowing for scalable approximate inference by leveraging a continuous-depth perspective [167]. Another line of work formulates UQ within stochastic optimal control problems, utilizing backward SDEs for gradient computation and uncertainty estimation [168]. Further research has focused on developing frameworks to explicitly quantify and propagate both aleatoric and epistemic uncertainties, such as using SDE-Recurrent Neural Networks for time-series data [169].
IV-D2 SDEs for Graphs
SDEs are increasingly employed with GNNs for UQ by modeling the evolution of graph representations as continuous-time stochastic processes. Bergna et al. propose the Latent Graph Neural Stochastic Differential Equations (LGNSDEs) framework to learn uncertainty-aware representations for graph data [76]. The core idea is to model the evolution of latent node features $\mathbf{Z}_{t}$ using an SDE. The framework defines an Ornstein-Uhlenbeck prior process and an approximate posterior SDE. The drift function $F_{G}$ of this posterior SDE is parameterized by a GCN, while the diffusion function $G_{G}$ is shared with the prior to ensure a bounded KL divergence. The approximate posterior SDE is given by:
$d\mathbf{Z}_{t}=F_{G}(\mathbf{Z}_{t},t,\boldsymbol{\phi})dt+G_{G}(\mathbf{Z}_{t},t)d\mathbf{W}_{t}$ (27)
where $\boldsymbol{\phi}$ represents the learned GCN weights and $d\mathbf{W}_{t}$ is a Wiener process.
GNSD [77] connects GNNs with Stochastic Partial Differential Equations (SPDEs). GNSD models an SPDE that ties the evolution of node representations to the graph structure. Drift and diffusion terms use graph operators such as Laplacians or neighborhood attention to drive node representation dynamics. For epistemic uncertainty, GNSD employs a specific stochastic forcing network and a Q-Wiener process defined on the graph domain to model its propagation among nodes, differing from LGNSDE’s Bayesian treatment of latent SDEs. In dynamic graph and generative modeling contexts, SDEs are utilized in decoders for inferring dynamic causal graphs in spatio-temporal forecasting [78] or to define continuous-time generative diffusion processes for permutation-invariant graph generation [79]. Furthermore, AGGDN model [80] stacks an SDE module upon an ODE component, with the ODE acting as a control signal to modulate the SDE propagation.
@@ V Methods for Uncertainty Handling
Effective handling is essential for developing reliable machine learning models. In this section, we explore four key approaches for managing uncertainty: direct inference, out-of-distribution techniques, calibration, and conformal prediction. These methods offer distinct strategies for addressing scenarios where models directly handle uncertainty, encounter data outside its training distribution, adjusting model confidence, and ensuring robust predictions.
V-A Bayesian Inference Techniques
In this section we introduce methods for uncertainty handling based on Bayesian inference. We first discuss exact Bayesian inference methods that compute posterior distributions directly from Bayes theorem. We then present approximate methods that use Monte Carlo dropout and Variational Inference for graph data.
V-A1 Exact Bayesian Inference
Exact Bayesian Inference Overview. Exact Bayesian inference refers to methods that derive closed-form posterior distributions by directly applying Eq. (14) to the parameters of interest. This approach naturally fits node classification tasks, as neighboring nodes can be assumed as observed data (or evidence) to infer a node’s posterior distribution of the parameters used for classification.
Exact Bayesian Inference for Graphs. SocNL [81] is one of the early works based on exact Bayesian inference. It modeled node labels as categorical random variables and used a Dirichlet prior to initialize label distributions. The method assumes that a node tends to share similar labels with its neighbors, thus using the neighbors’ labels as observed data to update the prior and derive the closed-form Dirichlet posterior.
Beyond the label-sharing assumption and the mere propagation of labels (i.e., directly using the neighbors’ labels as observed data), NETCONF [83] introduces the propagation of multinomial messages to support both homophily and heterophily network effects. It captures these effects by employing compatibility matrices to represent the strength of connections between nodes and then using the resulting counts as evidence in the closed-form update. To handle high-dimensional node features, GPN [82] first uses an MLP to encode each node’s features. Then, normalizing flows [170] are applied to compute class pseudo-counts as observed data. Finally, Personalized PageRank propagates these class pseudo-counts to update the Dirichlet prior parameters for each node. SocNL, GPN, and NETCONF all utilize Dirichlet priors and posteriors, with the key distinction lies in the choice of observed data.
V-A2 Approximate Bayesian Inference
Approximate Bayesian Inference Overview. Approximate Bayesian inference is a well-established tool for handling uncertainty when exact Bayesian methods are computationally intractable. In this section, we focus on two prominent techniques: Monte Carlo (MC) Dropout [171] and Variational Inference (VI) [172].
MC Dropout approximates the Bayesian posterior by interpreting the dropout mechanism as a variational approximation over the model weights. Specifically, given an input $\mathbf{x}$, the predictive distribution is estimated by averaging over multiple stochastic forward passes as follows:
$p(\mathbf{y}\mid\mathbf{x})\approx\frac{1}{T}\sum_{t=1}^{T}p\bigl{(}\mathbf{y}\mid\mathbf{x},\hat{\boldsymbol{\theta}}^{(t)}\bigr{)},$ (28)
where $\hat{\boldsymbol{\theta}}^{(t)}$ represents a sampled set of model weights induced by applying dropout at each forward pass, $T$ is the total number of Monte Carlo samples.
In contrast, VI seeks to approximate the true posterior $p(\boldsymbol{\theta}\mid\mathcal{D})$ by a tractable variational distribution $q_{\boldsymbol{\phi}}(\boldsymbol{\theta})$, parameterized by $\boldsymbol{\phi}$. This approximation is achieved by maximizing the Evidence Lower BOund (ELBO):
$\mathcal{L}(\boldsymbol{\phi})=\mathbb{E}_{q_{\boldsymbol{\phi}}(\boldsymbol{\theta})}\Bigl{[}\log p(\mathcal{D}\mid\boldsymbol{\theta})\Bigr{]}-\mathrm{KL}\Bigl{(}q_{\boldsymbol{\phi}}(\boldsymbol{\theta})\parallel p(\boldsymbol{\theta})\Bigr{)},$ (29)
which balances the data likelihood with a regularization term expressed as the Kullback–Leibler divergence between the variational distribution and the prior $p(\boldsymbol{\theta})$. VI is also a cornerstone in approaches such as VAEs [56], where latent variable modeling plays a central role.
Approximate Bayesian Inference for Graphs. Graph learning models often employ approximate Bayesian inference techniques to handle uncertainty over model parameters, latent representations, and even the graph structure itself. In many cases, these techniques follow the same principles as those outlined in Eqs. (28) and (29), but they are adapted to account for graph data.
Bayesian GCNNs [60] leverage MC dropout to approximate the posterior distribution given in Eq. (IV-A2). The predictive posterior of node is estimated as
$p(\mathbf{Y}^{*}\mid\mathbf{Y}_{\mathcal{L}},\mathbf{X},G_{obs})\approx\frac{1}{N_{G}S}\sum_{i=1}^{N_{G}}\sum_{s=1}^{S}p\bigl{(}\mathbf{Y}^{*}\mid\hat{\boldsymbol{\theta}}_{s,i},G_{i},\mathbf{X}\bigr{)},$ (30)
where $G_{i}$ denotes a graph sample drawn from an assortative mixed membership stochastic block model (e.g., an a-MMSBM [134, 135]), and $\hat{\boldsymbol{\theta}}_{s,i}$ is a stochastic sample of weights (obtained via MC dropout). Similarly, other works [120, 121, 62, 44, 76] also use MC dropout to approximate the predictive posterior in their graph learning models.
VI is also applied in graph settings to approximate the posterior over diverse latent variables. For instance, VGAE [58] obtains the solution for Eq. (16) by optimizing the ELBO:
$\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{Z}|\mathbf{X},\mathbf{A})}[\log\,p(\mathbf{A}|\mathbf{Z})]-\text{KL}\Bigl{[}q_{\boldsymbol{\phi}}(\mathbf{Z}|\mathbf{X},\mathbf{A})\,\|\,p(\mathbf{Z})\Bigr{]}.$ (31)
$\log p(\mathbf{A}|\mathbf{Z})$ represents the reconstruction term, which encourages the model to accurately reconstruct the graph’s adjacency matrix. The KL divergence term $\mathrm{KL}[q(\mathbf{Z}|\mathbf{X},\mathbf{A})\|p(\mathbf{Z})]$ acts as a regularizer, ensuring that the learned distribution does not deviate too far from the prior $p(\mathbf{Z})$. Other works that similarly use VI to estimate the predictive posterior include approaches for node representations [114, 115, 116, 117, 118, 119, 64, 65, 66, 123, 125, 126, 127, 76, 78], for graph representations [59], and for graph structure [63].
V-B Handling uncertainty due to OOD data
When a model encounters OOD data, its epistemic uncertainty on those data points is typically higher. To handle the uncertainty, a surge of research has focused on OOD generalization, and most of them propose novel network architectures and objective functions to improve the model performance on OOD test data. We organize the discussion and introduce the studies according to the taxonomy of OOD handling strategies [173], which categories the methods into distributionally robust optimization (DRO) [146, 147, 148, 149, 150, 67, 153, 154, 155, 156, 160, 161, 174, 84, 68, 85, 69], adversarial training (AT) [86, 87, 175, 71, 176, 70, 88, 72, 89, 90], and self-supervised learning (SSL) [177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 91, 74, 92, 75, 93, 187, 188].
V-B1 Distributionally Robust Optimization